If 20th term of H.P is 1 & 3th term is -1/17, then find it's largest term
 

Grade:11

1 Answers

Harshit Singh
3 years ago
Dear Student
given,
1/a+19d=1
a+19d=1.......(i)
1/(a+2d)= – 1/17
a+2d= – 17......(ii)
solving equation i and ii
d=18/17
a= – 325/17
from this we can easily find the required data

Thanks